OpenStudy (anonymous):

FAN & MEDAL!!! On the graph of the equation 4x + 5y = 20, what is the value of the y-intercept? -5 -4 4 5 I think its 4or 5 @AaronAndyson

OpenStudy (aaronandyson):

If you have 4x+5y = 20, for example, you can plug in the the value 0 for x and solve for y. This gives you the value of the for the y intercept: In your case, you would have 4(0) + 5y = 20 5y = 20 y = 4 y-intercept is (0, 4) then substitute y = 0 and solve for x: for 4x+5y = 20 you get 4x + 5(0) = 20 4x = 20 x = 5 x-intercept is (5, 0) Draw (0.4) and (5, 0) and a line between them. Use a ruler to make it look nice. It is easy when you get the hang of it.

OpenStudy (anonymous):

so I got it right I had put 4. but I was thinking 5 also but the answer is 4 right?

OpenStudy (aaronandyson):

@chrissy2430 you there?

OpenStudy (aaronandyson):

hello?

OpenStudy (anonymous):

yes sorry my computer is acting up. I have to click on something and click back on my post to see your response and to reply back to you. but im here

OpenStudy (anonymous):

@AaronAndyson you there :(

OpenStudy (aaronandyson):

yes four is correct :)

OpenStudy (anonymous):

can you check another for me please this one seems weird because I didn't get a negative

OpenStudy (aaronandyson):

but new post :)

OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 2) and (1, 5)? y-5=3(x-1) y-2=-3(x-0) ---> mine but ik there isn't a negative in this answer y+2=3(x+1) y+5=-3(x+1)

OpenStudy (aaronandyson):

is the slope 3?

OpenStudy (anonymous):

yes I got positive 3 but I did a new post for you to see I tagged you in it

OpenStudy (anonymous):

To find y intercept with equation in standard form, you simply have to find what value of y when multiplied by its coeefficient equals 20

OpenStudy (anonymous):

You do not need to find the slope or change to a different form of linear equation, those ways will give you the same answer but will take longer and introduce more chances for you to make a mistake @chrissy2430